Andretti Company has a single product called a Dak. The company normally produces and sells 88,000 Daks each year at a sellinComplete this question by entering your answers in the tabs below. Reg 1A Req 1B Reg 2 Reg 3 Req 4A to 4C Req 4D Reg 5 AssumeComplete this question by entering your answers in the tabs below. Reg 1A Reg 1B Reg 2 Req 3 Req 4A to 4C Reg 4D Reg 5 The coComplete this question by entering your answers in the tabs below. Reg 1A Reg 1B Reg 2 Reg 3 Req 4A to 4C Reg 4D Reg 5 Due toComplete this question by entering your answers in the tabs below. Req 1A Req 1B Reg 2 Reg 3 Req 4A to 4C Req 4D Req 5 Due toComplete this question by entering your answers in the tabs below. Reg 1A Req 1B Reg 2 Req3 Req 4A to 4C Reg 4D Reg 5 An outs
Answer #1

Answer to Question 1:

Question 1-a Calculation of incremental net operating income Incremental units: 114,400 - 88,000 = 26,400 units Incremental P

Answer to Question 2, 3, and 4-a:

Question 2 Calculate the break even price as follows: Sales - variable costs = Fixed costs (26,400 X sales price per unit) -

Answer to Question 4-b, 4-c and 4-d:

Question 4-b Total fixed manufacturing cost avoided = (792,000/12) X 2 months X 65% = 85,800 Total fixed selling cost avoided

Answer to Question 5:

Question 5 The relevant cost are those that can be avoided by purchasing from outside manufacturer. These costs are: Variable